Joanna, while driving 60 miles per hour gets to Birmingham 1 hour earlier then when she drives 50 miles per hour. How far does she drive?
Answer by mananth(16946)

Let the time taken at 50mph be x distance = 50x
time taken at 60 mph = x-1 hrs distance = 60(x-1)
The distance traveled is the same.

50x=60(x-1)
50x= 60x-60
x=6
she drives 300 miles.
M Ananth ----------mananth@hotmail.com